OpenStudy (anonymous):

When the quadratic equation y = (x − 3)(2x + 7) is written in standard form, what is the value of the coefficient "b"?

OpenStudy (anonymous):

y= (x-3)(2x+7) y=2x^2+x-21 y=ax^2+bx+c So b is????

OpenStudy (anonymous):

^what @MagicallyFarting said?

OpenStudy (anonymous):

2x^2+x-21 ax^2+bx+c b=1